OpenStudy (dls):

A right triangle has an area of 138 unit sq and perimeter of 60.94 units. Find the length of hypotenuse.

OpenStudy (dls):

@terenzreignz :P

terenzreignz (terenzreignz):

Always one for the weird problems... Let's try doing it algebraically... we need two values, the lengths of the legs of said right triangle... Let's call them x an dy

terenzreignz (terenzreignz):

x and y

terenzreignz (terenzreignz):

Now, the area of a right triangle is just half the product of the measures of the legs... so... \[\Large \frac12xy = 138\]

terenzreignz (terenzreignz):

And the perimeter of the right triangle is just the sum of the measures of the sides... so... \[\Large x+y+h=60.94\] where h is the length of the hypotenuse...

OpenStudy (dls):

yup,i wrote these 2 equations already ,now what :P

terenzreignz (terenzreignz):

Now, you solve the system? What else? :P

OpenStudy (dls):

lol

OpenStudy (dls):

x=60.94-y-h (60.94-y-h)(y)=138*2 o.O

terenzreignz (terenzreignz):

... \[\Large h = 60.94-x-y\]\[\Large \sqrt{x^2+y^2}=60.94-x-y\]

OpenStudy (dls):

yes,what after this?

terenzreignz (terenzreignz):

I don't know, square both sides? \[\Large x^2 + y^2 = (60.94-x-y)^2\]

OpenStudy (dls):

a-b-c omg

terenzreignz (terenzreignz):

huh?

OpenStudy (dls):

(a-b-c)^2 gotta be huge..hmm

terenzreignz (terenzreignz):

You're not exactly known for posting easy questions :P

terenzreignz (terenzreignz):

\[\Large x^2 + y ^2 = 60.94^2 -121.88(x+y)+(x+y)^2\]

OpenStudy (dls):

will we get answer from this? :P

terenzreignz (terenzreignz):

Let's see. Expand... \[\Large x^2 + y ^2 = 60.94^2 -121.88(x+y)+x^2+2xy + y^2\]\[\Large \cancel{x^2} + \cancel{y ^2} = 60.94^2 -121.88(x+y)+\cancel{x^2}+2xy + \cancel{y^2}\]

terenzreignz (terenzreignz):

\[\Large 0 = 60.94^2 -121.88(x+y) + 2xy\]

OpenStudy (dls):

and we know 2ab's value

OpenStudy (dls):

i mean 2xy

terenzreignz (terenzreignz):

Let's use this... \[\Large \frac12xy = 138\] Multiply both sides by 4, we get... \[\Large 2xy = 552\]

OpenStudy (dls):

yup

terenzreignz (terenzreignz):

\[\Large 0 = 60.94^2 -121.88(x+y)+552\]

OpenStudy (dls):

x+y is nearly 35

OpenStudy (dls):

(x+y)^2=x^2+y^2+2xy x^2+y^2=552

terenzreignz (terenzreignz):

Well, we could replace \(\Large y \) with \[\Large \frac{276}x\]

terenzreignz (terenzreignz):

wait what

OpenStudy (dls):

sqrt(552)=23.49 o.O

OpenStudy (dls):

??

terenzreignz (terenzreignz):

How?

OpenStudy (dls):

I got (x+y) from ur equation i.e 35 now expand \[\Huge (x+y)^2=x^2+y^2+2xy\] now isolate x^2+y^2 and take the sqrt

terenzreignz (terenzreignz):

I see. Well, good, I suppose, did you really need help with this? :P

terenzreignz (terenzreignz):

I'm not good at spotting the quick ways to solve problems... better fix that :D

OpenStudy (dls):

we are getting sqrt{552} it should be sqrt{673} hmm..

OpenStudy (dls):

okay i got it :P

terenzreignz (terenzreignz):

Great... LOL

terenzreignz (terenzreignz):

So... why did you drag me into it? :P

OpenStudy (dls):

to spend some time :D

terenzreignz (terenzreignz):

Well, time well-spent, I suppose :P

OpenStudy (dls):

haha